What is the pH of a 0.250 M solution of anilinium chloride, C6H5NH3CI? For aniline, C&HNH2, Kb = 4.2 x 10-10 C al ancwors
Answer #1

It has hydrolysis:

A- + H2O = HA + OH-

From the expression of Kb we have:

Kb = [HA] * [OH-] / [A-] = X ^ 2 / 0.25 - X

It is assumed that - X is negligible and clears:

X = [OH-] = √Kb * 0.25 = √4.2x10 ^ -10 * 0.25 = 1.02x10 ^ -5 M

The pOH and the pH are calculated:

pOH = - log 1.02x10 ^ -5 = 5

pH = 14 - 5 = 9

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.